OpenStudy (anonymous):

elimination; 6x - 3y = 15 7x + 4y = 10

OpenStudy (anonymous):

(6x-3y=15)4 (7x+4y=10)3

OpenStudy (anonymous):

24x-12y=60 21x+12y=30

OpenStudy (anonymous):

45x=90

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

x=2 y=1